Status Report

NASA ISS On-Orbit Status 3 May 2008

By SpaceRef Editor
May 3, 2008
Filed under , , ,
NASA ISS On-Orbit Status 3 May 2008

All ISS systems continue to function nominally, except those noted previously or below. Saturday – half-day off for CDR Volkov, FE-1 Kononenko & FE-2 Reisman.

To provide cooling for the ground-commanded activation of the U.S. CDRA (Carbon Dioxide Removal Assembly), FE-2 Reisman begun his workday by connecting the regular ITCS LTL (Internal Thermal Control System/Low Temperature Loop) coolant jumper connection to the LAB1D6 rack. [CDRA activation took place at 4:00am-5:15am EDT. The CDRA will operate over the weekend. Deactivation will occur when ppCO2 drops to 2.5 mmHG. Deactivation will be Sunday evening or early Monday morning.]

The crew conducted the regular weekly three-hour task of thorough house cleaning. ["Uborka", normally done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab. Additionally, as part of the weekly cleaning, Malenchenko performs an inspection of structural elements, cables and instruments behind SM panels for moisture.]

As part of today’s uborka, Volkov & Kononenko conducted preventive maintenance cleaning on the FS5, FS6, FS9, VPkhO, VdPrK, & VPrK fan grilles in the SM and the TsV2 fan inlet in the FGB (Funktsionalnyi-Grusovoi Blok, Functional Cargo Block).

At ~10:00am, the crew held their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

Sergey Volkov conducted the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables. [Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, replacement of the KOV EDV for the Elektron-intended water, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.]

The crew completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exerciser (FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1). [The RED, as an anaerobic muscle exerciser, allows a variety of routines: squat, heel raises, bent-over rowing, abdominal crunches, deadlift, bench presses, upright rowing, etc. For Sergey & Oleg, who are using RED three times a week, each session features four different routines which vary from day to day to target different muscle groups.]

Afterwards, Garrett transferred the crew’s exercise data file to the MEC (Medical Equipment Computer) laptop for downlinking, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Working off the discretionary “time permitting” task list, Volkov completed another KPT-3 session to make observations and take aerial KPT-3 photography of environmental conditions for Russia’s Environmental Safety Agency (ECON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens.

Also from the suggestions list, Kononenko performed a session of the Russian GFI-8 "Uragan" (hurricane) earth-imaging program, using the NIKON D2X to take telephotos of the hydrological situation on the Volga Aktyubinsk alluvial plain and oil spills on the Caspian Sea.

Weekly Science Update (Expedition Seventeen — Week 2)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module. Inc16 science data returned on 15S, waiting for hand-over to science team.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Germination in B3 and B4. A1 plants are still alive. Investigators are evaluating the possibility of EC A1/B1 position change.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): The final run of the CSLM-2 SPU was completed. The data files have been downlinked and the temperature data looks good. Everything appears to have worked as expected. Detailed analysis of the data will be forth coming as well as inspection of the samples once they return to the ground. “Thank you for supporting CSLM-2 research!!”

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): The crew was thanked for supporting the ELITE experiment. The data files were successfully downlinked and data analysis on the ground can now be started.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed. Inc16 science data returned on 15S, waiting for hand-over to science team.

EuTEF (European Technology Exposure Facility): On 4/25, EuTEF platform encountered an accidental de-activation of its power feeder #1. Ground operators re-activated all the instruments progressively, with the exception of FIPEX, which got delayed by a couple of days. – DEBIE-2: Link error still under investigation; – DOSTEL: On-going science acquisition; – EuTEMP: Currently inactive as planned; – EVC: Further troubleshooting is on-going; still problems to get high-rate data downlink; – EXPOSE: On-going science acquisition; – FIPEX: Long-duration measurements with sensors RAM1, RAM4 and ZENITH8 has been abruptly interrupted on 4/25 after the accidental power feeder #1 de-activation. New science acquisition script has been started on 4/30; – MEDET: On-going science acquisition, running nominal; – PLEGPAY: Instrument is powered on, but not in science acquisition mode; – TRIBOLAB: On 4/25, the very sensitive TRIBOLAB Pin On Disk #2 (POD #2) run could be restarted, but science impact assessment is on-going.

FSL (Fluid Science Laboratory): FSL LAN (Local Area Network) was successfully performed on 4/30. Further troubleshooting activities are required after the erroneous telemetry with CEM-Upper) Optical Module. FSL MIL Bus cable has still to be repaired.

GEOFLOW: Start of GEOFLOW is pending further FSL commissioning activities (check of FSL optical modes) currently planned for next two weeks.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Garrett, your Increment 17 remaining Sleep activities include: 2 download/initialization sessions, 1 more week of required Sleep logging, and doffing the Actiwatch on the Shuttle on your return home. Any additional sleep logging than the required week is above and beyond and greatly appreciated by the PI. Also, feel free to get a new Actiwatch band from the Actiwatch Supply Kit (#1042) and discard your old one. Just call down so we can track it.”

SOLAR (Solar Monitoring Observatory): Initial Sun tracking was started on 4/29, but the Sun was not in the proper beta range to allow for nominal data acquisition – SOLAR platform went to Stand-By Mode when trying to track the Sun. On 4/30, solar measurements started with SOVIM and SOLSPEC instruments, but were affected by short Sun passes. – SOVIM: Acquiring science data – some glitches observed in the instrument telemetry; – SOLSPEC: Calibration activities (3 out of 4) successfully performed on 5/01. Acquiring science data; – SOLACES: Science acquisition to be restarted on 5/02, after some Command Sequence script was stopped on 5/01. Planned measurements will complement the ones gathered during the NASA Sounding Rocket campaign last April.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The 4 ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B.

CEO (Crew Earth Observations): Through 4/30 the ground has received a total of 560 ISS/CEO images from Inc17 for review and cataloging. Currently under review are images with camera times corresponding to the following target list requests: Harvard Forest, Vermont (target acquisition unverified yet); Tigris-Euphrates Delta (preliminarily-appears somewhat overexposed); Tropical Cyclone Nargis, Bay of Bengal (good effort); Afar Rift Zone, Ethiopia (preliminarily – very good with much of the area acquired); and Lake Pontchartrain, Mississippi Delta (excellent coverage of this dynamic event – will be published). “We are attempting to respond to your request for support in acquiring good views of the cities of Montreal, Cologne, Philadelphia, and Washington D.C. We have other targets in these areas as well, but have not been able to ask for them either because of poor weather. Now it appears we will have to wait a few weeks before illumination become adequate again for these locations. We will report back if we find any views of them in your imagery. Most of your imagery appears to be focused and well-composed, however please be advised that: a small number of your practice or personal interest images using the long lens settings and having no context views are extremely difficult to catalog and may never be located precisely. Thank you for your great response to our target requests to date!”

CEO (Crew Earth Observations) photo targets uplinked for today were Tigris-Euphrates Delta, Middle East (ISS orbit track brought the station directly over the historically significant region that feeds into the Persian Gulf to the SE. Overlapping frames, taken along track, will provide a useful contextual view of the current configuration of the delta), Ile Rouleau Impact Crater, Quebec, Canada (weather was predicted to be clear over Ile Rouleau, a 4 km-diameter impact structure that is exposed at the surface as a small island. Looking to the right of track for the long, narrow northeast-southwest oriented Lake Mistassini – in appearance, the lake looks somewhat like claw marks on the landscape; overlapping frames along the extent of the lake will capture the crater near the southern end), Montreal, Quebec, Canada (Crew Request. Looking to the right of track for Montreal, weather was predicted to be clear), Santa Barbara Coast, California (ISS had a nadir pass over this long-term ecological research [LTER] site. Overlapping frames, taken along track as approaching the coast, were requested. Continuing to acquire overlapping frames as ISS crossed the urban area of Santa Barbara and crossed the coastline once again), and Madrean Sky Islands, Mexico (looking to the left of track as the station traversed the Gulf of California towards the Sierra Madre in Mexico. Contextual frames of the mountains, taken as ISS traveled parallel to them from NW to SE, will be useful for assessing the status of the forested "sky islands" on the upper slopes).

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

ISS Orbit (as of this morning, 4:34am EDT [= epoch]):
Mean altitude — 341.8 km
Apogee height — 346.6 km
Perigee height — 337.0 km
Period — 91.37 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0007109
Solar Beta Angle — -10.4 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.76
Mean altitude loss in the last 24 hours — 97 m
Revolutions since FGB/Zarya launch (Nov. 98) — 54145

Significant Events Ahead (all dates Eastern Time, some changes possible.):
05/14/08 — Progress M-64/29P launch
05/16/08 — Progress M-64/29P docking (FGB nadir)
05/31/08 — STS-124/Discovery/1J launch – JEM PM “Kibo”, racks, RMS (5:01pm EDT)
06/02/08 — STS-124/Discovery/1J docking
07/10/08 — Russian EVA-20 (7/10-11)
08/10/08 — ATV1 undocking
08/12/08 — Progress M-65/30P launch
08/14/08 — Progress M-65/30P docking (SM aft port)
????/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
09/09/08 — Progress M-64/29P undocking (from DC1)
09/10/08 — Progress M-66/31P launch
09/12/08 — Progress M-66/31P docking (DC1)
10/01/08 — NASA 50 Years
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (SM aft port)
10/16/08 — STS-126/Discovery/ULF2 launch – MPLM Leonardo, LMC
10/18/08 — STS-126/Discovery/ULF2 docking
10/23/08 — Soyuz TMA-12/16S undocking (FGB nadir)
11/03/08 — Soyuz TMA-13/17S relocation (from SM aft to FGB nadir)
11/20/08 — ISS 10 Years
11/26/08 — Progress M-67/32P launch
11/28/08 — Progress M-67/32P docking (SM aft port)
12/04/08 — STS-119/Discovery/15A launch – S6 truss segment
12/06/08 — STS-119/Discovery/15A docking
12/15/08 — STS-119/Discovery/15A undocking
2QTR CY09 — STS-127/2J/A launch – JEM EF, ELM-ES, ICC-VLD
3QTR CY09 — STS-128/17A/Atlantis – MPLM(P), last crew rotation
05/??/09 — Six-person crew on ISS (following Soyuz 18S-2 docking)
3QTR CY09 — STS-129/ULF3/Discovery – ELC1, ELC2
4QTR CY09 — STS-130/20A/Endeavour – Node-3 + Cupola
1QTR CY10 — STS-131/19A/Atlantis – MPLM(P)
1QTR CY10 — STS-132/ULF4/Discovery – ICC-VLD, MRM1 (contingency)
2QTR CY10 — STS-133/ULF5/Endeavour – ELC3, ELC4 (contingency).

SpaceRef staff editor.